Question 14 z Company uses the formula y-a + bX to predict and analyze overhead costs. In the previous year, XYZ used $1,750 per month for the a factor and $0.35 for the b factor in applying overhead. XYZ has used direct labour hours in the past, but is wondering whether overhead behaviour is more closely associated with machine hours. The following data have been generated for Month Machine Hours Overhead Costs 432 468 394 483 $2,646 3,089 2,476 $2,903 $2,710 418 Determine the fixed and variable values, using machine hours. (Round variable cost to 2 decimal places,e-g. 15.25 and fixed cost to 0 decimal places, e.g. 525) Fixed Cost Variable Cost Round answer to o decimal places, e.g. 525). it is expected that the company will Incur 896 machine hours in the coming month, predict that months overhead costs. (
BACK Ih lt Is expected that the company will incur 896 machine hours in the coming month, predict that months overhead costs. (Round answer to O decimal places, e.g. 525) Months overhead costs For April of the coming year, XYZ expects 3,265 direct labour hours and 679 machine hours. Predict overhead costs using (1) direct labour hours. (Round answer to O decimal place, e.g. 1,52s.) Overhead cost (2) machine hours. (Round answer to o decimal place, e.g. 1,525.) Overhead cost s NK TO TEXT If Aprils actual fixed overhead costs are $1,674 and actual variable overhead costs are $1,392, which activity base appears better? Better activity base
Answer #1

a) Determine fixed and variable values

Variable values = Change in Cost/Change in machine hour = (2710-2476)/(509-394) = 2.03 per machine hour

Fixed Cost = 2710-(509*2.03) = $1677

b) Month's overhead cost = 1677+(896*2.03) = $3496

1) Direct labor hours = 1750+(3265*.35) = 2893

2) Machine hour = 1677+(679*2.03) = $3055

3) Better activity base = Machine hour